HENDI SPECTROSCOPY: A REVIEW OF RECENT RESULTS

\underline{G. SCOLES}, Department of Chemistry, Princeton University, Princeton, N.J. 08544.

The recently established research field of Helium Nano Droplet Isolation (HENDI) spectroscopy will be reviewed with an emphasis on two of its most active areas: a) Use of HENDI for the formation and study of highly non-equilibrium aggregates of atoms and molecules and b) The problem of the measurement and calculations of the contribution of the quantum matrix (He and/or H_{2}) to the inertia of the molecules imbedded in it. This work was carried out with the support of the AFOSR and the NSF and with the collaboration of K. K. Lehmann, C. Callegari, A. Conjusteau, R. Schmied, and M. Radcliff.